# Correlators are simpler than wavefunctions

### v2

@@ Appendix A Some Derivations
$\mathcal{I}=\sum_{p_{*}\in\{x_{*},x,y_{e}\}}\mathop{\mathrm{Res}}_{p\to ip_{*}}\frac{2x_{*}x\prod_{e\in E(P_{*})}2y_{e}}{(p^{2}+x_{*}^{2})(p^{2}+x^{2})\prod_{e\in E(P_{*})}(p^{2}+y_{e}^{2})}$ (45)
@@
$\langle G\rangle^{(1)}=\epsilon\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}ds^{E}F(0)\Bigg\{\left(\sum_{v}\bar{x}_{v}S_{v}\right)^{2}-2\sum_{v}\bar{x}_{v}S_{v}^{2}\delta(\tau)\Bigg\}\prod_{e\in E(G)}G_{F}(s_{e}),$ (46)
@@
Recall that $S_{v}=\sum_{e\in P_{*}^{v}}s_{e}$, and $P_{*}^{v}$ is the set of edges that connect the vertex $v$ to the root vertex. It is clear that all cross terms in the square (like $s_{e}s_{e^{\prime}}$) yield odd integrands, which vanish when integrated. Then, it’s important to consider that for a given edge, $e$, we can split the graph in two connected sub-trees. One will have the root vertex, and the other will not. Then, for the edge $e$, will call the set of vertices that are in the connected sub-tree which does not contain the root vertex to be $V_{*}^{e}$, and $\bar{V}_{*}^{e}$ to be it’s complement (the vertices on the sub-tree with the root vertex). Now we can revert the sums over vertex variables, $x_{v}$, and integration variables $s_{e}$, to yield:
$\sum_{e\in E}s_{e}^{2}\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)^{2}-\left(\sum_{i}^{V}x_{i}\right)\left(\sum_{e\in E}s_{e}^{2}\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)=\sum_{e\in E}s_{e}^{2}\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)^{2}-\sum_{e\in E}s_{e}^{2}\left(\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)^{2}+\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)\left(\sum_{j\in\bar{V}_{*}^{e}}\bar{x}_{j}\right)\right)\,.$ (48)
There are three terms on the RHS, but the first and second cancel and we are left with:
$\sum_{e\in E}s_{e}^{2}\left(\sum_{u\in V_{*}^{e}}\sum_{v\in\bar{V}_{*}^{e}}\bar{x}_{u}\bar{x}_{v}\right)=-\sum_{u\neq v}x_{u}x_{v}\left(\sum_{e\in P_{u}^{v}}s_{e}^{2}\right)\,$, (49)
where we have introduced the notation $P_{u}^{v}$ for the path between $u$ to $v$. The jump from the LHS to the RHS can be seen by considering which edge integration variables $s_{e}$ are multiplying a given term $x_{u}x_{v}$. If we go back to consider the two sub-trees in which the graph splits with respect to the edge $e$, then a given $s_{e}$ in the sum can only multiply to vertex variables if the respective vertices sit on opposite sub-trees. Thus, only edges on the path from $u$ to $v$ can enter the sum on the RHS. So we are finally left with
$\langle\widehat{G}\rangle^{(1)}=\frac{-\epsilon}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\int_{-\infty}^{\infty}ds^{E}\sum_{e\in P_{u}^{v}}s_{e}^{2}\prod_{vv^{\prime}\in E_{G}}\widehat{G}_{F}(s_{e}),$ (50)

### v3

@@ Appendix A Some Derivations
$\mathcal{I}=2i\sum_{p_{*}\in\{x_{*},x,y_{e}\}}\mathop{\mathrm{Res}}_{p\to ip_{*}}\frac{2x_{*}x\prod_{e\in E(P_{*})}2y_{e}}{(p^{2}+x_{*}^{2})(p^{2}+x^{2})\prod_{e\in E(P_{*})}(p^{2}+y_{e}^{2})}$ (45)
@@
$\langle G\rangle^{(1)}=\frac{\epsilon}{2}\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}ds^{E}F(0)\Bigg\{\left(\sum_{v}\bar{x}_{v}S_{v}\right)^{2}-2\sum_{v}\bar{x}_{v}S_{v}^{2}\delta(\tau)\Bigg\}\prod_{e\in E(G)}G_{F}(s_{e}),$ (46)
@@
The expression in curly brackets becomes:
$\left(\sum_{v}\bar{x}_{v}S_{v}\right)^{2}-\sum_{u,v}\bar{x}_{u}\bar{x}_{v}S_{v}^{2}=-\frac{1}{2}\sum_{u,v}\bar{x}_{u}\bar{x}_{v}\left(S_{u}-S_{v}\right)^{2}=-\frac{1}{2}\sum_{u\neq v}\bar{x}_{u}\bar{x}_{v}\left(\sum_{e\in P_{u}^{v}}s_{e}\right)^{2}$ (48)
where we have introduced the notation $P_{u}^{v}$ for the path between $u$ to $v$. Thus, only edges on the path from $u$ to $v$ can enter the sum on the RHS. Considering that the cross terms vanish upon integration, we are finally left with
$\langle\widehat{G}\rangle^{(1)}=-\frac{1}{2}\frac{\epsilon}{\sum_{v}\bar{x}_{v}}\sum_{u\neq v}\bar{x}_{u}\bar{x}_{v}\int_{-\infty}^{\infty}ds^{E}\sum_{e\in P_{u}^{v}}s_{e}^{2}\prod_{vv^{\prime}\in E_{G}}\widehat{G}_{F}(s_{e}),$ (49)